# Harmonizing Generalization and Specialization: Uncertainty-Informed Collaborative Learning for Semi-supervised Medical Image Segmentation

### v1

@@ 4.1.1 Datasets
We evaluate UnCoL on three publicly available medical image segmentation benchmarks, none of which overlap with the pretraining data of the foundation models: (1) OASIS (2D Brain MRIs) [35]: This dataset comprises T1-weighted brain MRIs from 414 subjects. We perform five independent trials with a 7:1:2 split for training, validation, and test, respectively. The task is four-class tissue segmentation: Left/Right Thalamus (LT, RT), Hippocampus (LH, RH). All slices are resized to with an in-plane spacing of 1.0 and augmented with random rotation and flipping, following [15, 13]. (2) Pancreas-CT (3D Abdominal CT Images) [36]: This dataset contains 82 abdominal CT volumes with manual pancreas annotations. We follow the widely adopted semi-supervised setting in [22, 20, 37], using 62 volumes for training and 20 for testing, with the labeled subset identical to that used in prior works to ensure fair comparison. (3) ImageTBAD (3D Aortic CTA Scans) [38]: This dataset contains 100 thoracic CTA scans annotated for type B aortic dissection. We follow the same 7:1:2 train/validation/test split protocol, conducted at the volume level across five independent trials. To ensure compatibility with the input size of the foundation models, all volumes are first resampled to an isotropic resolution of 1.0 and then uniformly resized to a fixed shape of voxels using trilinear interpolation. Results are reported as the mean and standard deviation across trials.
@@ 4.1.2 Model Setup
We employ MedSAM [6] and SAM-Med3D [7] as the frozen 2D and 3D foundation models, respectively. Both models use a 12-layer Vision Transformer (ViT) as the image encoder, with a hidden dimension of 768. The student model employs a lightweight 4-layer SimpleViT encoder, paired with a task-specific backbone: U-Net [31] for 2D and V-Net [33] for 3D segmentation. Cross-layer distillation is conducted between the 4-layer transformer encoder of the student model and the 12-layer image encoder of the generalized teacher. Specifically, the first three student layers are aligned with the 4th, 8th, and 12th layers of the teacher encoder, a mapping selected to capture progressively higher levels of semantic priors. The final student layer is supervised by the output of the teacher’s image–prompt feature fusion module.
@@ 4.1.3 Training Details
All models are implemented in PyTorch and trained on a single NVIDIA RTX 3090 GPU. Training is performed using SGD with a learning rate of 0.01 and a weight decay of , for 15,000 iterations in both the pretraining and fine-tuning stages. During pretraining (Stage 1), each training batch consists of 4 labeled samples, while in fine-tuning (Stage 2), each batch includes 2 labeled and 2 unlabeled samples. Throughout training, synthetic prompts are supplied to the frozen generalized teacher model to guide supervision. Prompts are used only to query the frozen teacher, while the student model always performs prompt-free inference. For 2D inputs, bounding-box prompts are created by shifting ground-truth annotations or pseudo-label masks by 20 pixels. For 3D inputs, point prompts are sampled either from labeled voxels or from regions with low predictive uncertainty. Following [3], we adopt a Gaussian ramp-up function , where denotes the current iteration and , to progressively increase the supervision strength. The distillation weight is set as with , while the entropy threshold is defined as to gradually relax the filtering criterion [4]. Only voxels with entropy below are supervised, with a fixed loss weight [13].
@@ 4.2.1 Comparison on 2D MRI Segmentation (OASIS)
We evaluate UnCoL on the OASIS dataset across different annotation ratios. As shown in Table 1, UnCoL consistently outperforms existing methods using only 5–10% labeled data. While MedSAM [6], pretrained on brain MRI datasets like FeTA [39], achieves reasonable zero-shot performance (average Dice score of 76.49%), its effectiveness is limited by anatomical and distributional shifts, reflecting the general limitations of zero-shot foundation models under domain shift. UnCoL overcomes this by using uncertainty-guided supervision to adapt effectively with minimal labels.
@@
We observe two major failure modes in prior methods in Fig. 5: (1) TL–FL boundary leakage in the dissection plane due to low contrast, and (2) misclassification of thrombosed FL as background or TL, especially near the entry tear. Semi-supervised baselines often amplify these errors because pseudo labels are unreliable in such ambiguous regions. UnCoL mitigates these issues by combining global vascular priors from the foundation model with localized refinements from the specialized teacher, enabling uncertainty-guided corrections particularly beneficial for the high-uncertainty FL regions.
| Method | Model | DPKD | UAPL | OASIS | Pancreas-CT | ImageTBAD | ||||||||||||||
| Stu. | G-Tch | S-Tch | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | ||||||
| Baseline | ✓ | 66.395.55 | 56.715.04 | 24.011.55 | 9.843.17 | 56.9616.50 | 41.3813.66 | 18.6912.60 | 5.383.90 | 55.8521.20 | 42.4118.96 | 14.1911.74 | 3.654.56 | |||||||
| ✓ | 76.490.67 | 62.680.65 | 7.300.07 | 2.230.04 | 79.374.55 | 66.026.22 | 5.521.79 | 1.550.54 | 29.61 19.11 | 19.04 14.43 | 13.23 5.58 | 4.44 1.94 | ||||||||
| ✓ | ✓ | 84.024.15 | 76.413.80 | 12.863.58 | 4.371.72 | 64.2811.62 | 48.4012.10 | 28.1914.54 | 8.925.15 | 64.6715.93 | 50.0415.83 | 7.486.06 | 2.001.43 | |||||||
✓ | ✓ | ✓ | 93.652.54 | 88.164.26 | 3.149.13 | 0.901.62 | 73.229.95 | 58.6811.44 | 10.396.57 | 3.381.52 | 52.8316.38 | 41.2716.88 | 4.434.75 | 1.301.27 | |||||
| ✓ | ✓ | ✓ | 93.172.17 | 87.293.71 | 2.395.52 | 0.831.28 | 72.669.96 | 57.9811.68 | 11.905.92 | 3.761.60 | 67.6316.48 | 53.3317.06 | 5.444.84 | 1.631.60 | ||||||
| ✓ | ✓ | ✓ | ✓ | 93.693.75 | 88.346.03 | 1.950.94 | 0.610.50 | 74.129.49 | 59.7411.09 | 10.796.63 | 3.381.78 | 69.7317.56 | 56.0218.65 | 5.927.43 | 1.641.38 | |||||
✓ | ✓ | ✓ | ✓ | ✓ | S-Tch-PL | 93.082.32 | 87.143.91 | 2.445.28 | 0.770.87 | 80.086.11 | 67.188.05 | 7.204.10 | 2.151.08 | 71.261.58 | 57.441.54 | 4.841.75 | 1.380.40 | ||
| ✓ | ✓ | ✓ | ✓ | ✓ | Avg-Tch-PL | 92.972.32 | 86.953.90 | 2.485.24 | 0.760.77 | 77.297.58 | 63.569.20 | 7.575.13 | 2.321.40 | 70.791.68 | 57.161.22 | 4.871.77 | 1.300.40 | |||
| ✓ | ✓ | ✓ | ✓ | ✓ | UAPL | 94.420.29 | 89.510.47 | 2.063.27 | 0.610.63 | 80.405.82 | 67.607.83 | 6.863.96 | 1.850.94 | 70.976.35 | 57.176.81 | 4.832.22 | 1.300.58 | |||
@@ 4.3.1 Effect of Teacher Model Configuration
To assess the influence of different teacher configurations, we first compare the baseline model without teacher guidance, single-teacher settings, and the dual-teacher setup. As shown in the top block of Table 4, the Generalized Teacher (G-Tch) increases Dice over the student-only baseline from 66.39% to 76.49% on OASIS and from 56.96% to 79.37% on Pancreas-CT, demonstrating effective prior transfer from the foundation model. However, when encountering unseen anatomy, its performance collapses to 29.61% on ImageTBAD (vs. 55.85% for the student), even with ground-truth–derived prompts, suggesting its domain dependency. The Specialized Teacher (S-Tch), derived via an EMA of the student model, provides stable gains (84.02%, 64.28%, and 64.67%) but underperforms the generalized teacher on well-studied domains (e.g., 64.28% vs. 79.37% on Pancreas-CT). These results show that neither teacher alone is consistently reliable, motivating a dual-teacher design to balance generalization and specialization. Since directly combining two imperfect teachers can propagate errors, DPKD and UAPL are introduced to selectively transfer reliable knowledge, as validated in later ablations.
@@ 4.3.3 Pseudo-Label Quality Analysis
As shown in Table 5, the S-Tch produces more accurate and stable pseudo-labels than the G-Tch, achieving Dice scores of 94.35%, 72.77%, and 66.48% on OASIS, Pancreas-CT, and ImageTBAD, respectively. This improvement is consistent with the effect of transferring generalized semantics, as evidenced by the S-Tch baseline without G-Tch assistance in Table 4 (+10.33%, +8.49%, and +1.61% Dice gains). The G-Tch performs well on OASIS (82.11%) and Pancreas-CT (60.24%) but drops to 15.98% on ImageTBAD, indicating limited domain adaptability. Meanwhile, the student achieves comparable or slightly higher accuracy (e.g., +3.25% Dice on ImageTBAD), demonstrating effective knowledge integration from both teachers and resilience to negative transfer. Overall, these results suggest that our pretraining yields reliable pseudo-labels that adapt well to dataset-specific characteristics, providing a solid foundation for uncertainty-aware fine-tuning.
@@ 4.3.5 Effect of on Model Performance
We evaluate the sensitivity of UnCoL to the pseudo-label balancing coefficient , which controls the trade-off between supervision from ground-truth labels and pseudo-labels during semi-supervised training. Consistent with prior work [13], we set as the default. Figure 7 shows that UnCoL maintains strong performance across a wide range of values. On ImageTBAD, the highest Dice score of 79.24% is obtained at , while Pancreas-CT reaches its peak Dice of 83.91% at . Boundary-based metrics such as the ASD exhibit similar stability. These results suggest that UnCoL is relatively insensitive to the choice of , with the uncertainty-aware pseudo-labeling mechanism consistently improving performance without the need for precise hyperparameter tuning.
@@ 4.4 Comparison with Classic Fine-tuning Strategies
Experiments in Table 6 follow the fine-tuning paradigm summarized in [41], which provides a unified analysis of classical adaptation strategies for medical foundation models. Our method demonstrates two practical advantages over these tuning baselines. First, even with only limited labeled data, our framework consistently surpasses the fully supervised fine-tuned models, suggesting enhanced label efficiency and improved adaptation under scarce annotations. Notably, across complex datasets such as OASIS and ImageTBAD, fine-tuning with limited data offers little improvement over the zero-shot baseline, underscoring the difficulty of adapting large medical foundation models when annotations are scarce. Furthermore, on ImageTBAD, even full fine-tuning does not surpass the un-tuned model, revealing that parameter adaptation remains ineffective under severe distributional shifts. Second, our method operates in a fully prompt-free manner, whereas the SAM-based fine-tuning baselines require test-time prompts (e.g., bounding boxes or points), introducing additional annotation overhead. These results highlight that our method attains superior label efficiency and more robust adaptation to complex anatomical structures.
@@ 4.5.1 Discriminability
The discriminative power of uncertainty is measured using the area under the ROC curve (AUROC), which quantifies how well uncertainty distinguishes between correct and incorrect predictions. Higher AUROC indicates stronger correlation between uncertainty and voxel-wise error. As shown in Fig. 8(a), UnCoL achieves the highest AUROC on OASIS (0.995) and Pancreas-CT (0.982), and performs competitively on ImageTBAD (0.939). The relatively lower AUROC on ImageTBAD is consistent with the presence of domain shifts and complex anatomical variations. Nonetheless, UnCoL still achieves the best Dice and boundary metrics on this dataset (Table 3), indicating that uncertainty estimation remains reliable even in challenging settings. Overall, these results suggest that our dual-teacher framework improves both segmentation accuracy and the discriminative capacity of uncertainty.
@@ 4.5.2 Calibration
Reliability diagrams provide a standard visualization of the relationship between predictive confidence and observed accuracy [42]. Perfect calibration corresponds to points lying on the diagonal (confidence = accuracy). As shown in Fig. 8(b), UnCoL exhibits strong calibration across all datasets, with small confidence deviations and low calibration errors measured by Expected Calibration Error (ECE ), Negative Log-Likelihood (NLL ), and Brier score (). These quantitative measures confirm that UnCoL’s predicted confidence values are well aligned with voxel-level correctness.
@@ 4.5.3 Selective prediction reliability
Risk–coverage curves evaluate how uncertainty supports selective prediction by progressively excluding high-uncertainty voxels [43]. A reliable uncertainty measure yields lower selective risk at any target coverage. As shown in Fig. 8(c), UnCoL’s entropy-based uncertainty consistently achieves the lowest risk compared with energy scores, MC Dropout variance, and MC Dropout–based mutual information, indicating that regions assessed as low-uncertainty tend to correspond to more reliable predictions.
Collectively, these results indicate that UnCoL produces uncertainty estimates that are both discriminative and well-calibrated, effectively supporting pseudo-label selection and consistency regularization under domain-shifted and data-scarce scenarios.

### v2

@@ 4.1.1 Datasets
We evaluate UnCoL on three publicly available medical image segmentation benchmarks that are not included in the pretraining datasets of the foundation models. (1) OASIS (2D Brain MRIs) [26]: This dataset comprises T1-weighted brain MRIs from 414 subjects. We perform five independent trials with a 7:1:2 split for training, validation, and test, respectively. The task is four-class tissue segmentation: Left/Right Thalamus (LT, RT), Hippocampus (LH, RH). All slices are resized to with an in-plane spacing of 1.0 and augmented with random rotation and flipping, following [22, 3]. (2) Pancreas-CT (3D Abdominal CT Images) [31]: This dataset contains 82 abdominal CT volumes with manual pancreas annotations. We follow the widely adopted semi-supervised setting in [37, 38, 21], using 62 volumes for training and 20 for testing, with the labeled subset identical to that used in prior works to ensure fair comparison. (3) ImageTBAD (3D Aortic CTA Scans) [40]: This dataset contains 100 thoracic CTA scans annotated for type B aortic dissection. We follow same 7:1:2 train/validation/test split at the volume level across five trials. To ensure compatibility with the input size of the foundation models, all volumes are first resampled to an isotropic resolution of 1.0 and then uniformly resized to a fixed shape of voxels using trilinear interpolation. Results are reported as mean and standard deviation across trials.
@@ 4.1.2 Model Setup
We employ MedSAM [23] and SAM-Med3D [34] as the frozen 2D and 3D foundation models, respectively. Both models use a 12-layer Vision Transformer (ViT) as the image encoder, with a hidden dimension of 768. The student model employs a lightweight 4-layer SimpleViT encoder, with a task-specific backbone: U-Net [30] for 2D and V-Net [29] for 3D segmentation. Cross-layer distillation is conducted between the 4-layer transformer encoder of the student model and the 12-layer image encoder of the generalized teacher. Specifically, the first three student layers are aligned with the 4th, 8th, and 12th layers of the teacher encoder, to capture progressively higher levels of semantic priors. The final student layer is supervised by the output of the teacher’s image-prompt fusion module.
@@ 4.1.3 Training Details
All models are implemented in PyTorch and trained on a single NVIDIA RTX 3090 GPU. Training uses SGD with a learning rate of 0.01 and a weight decay of , for 15,000 iterations in pretraining and fine-tuning stages. During pretraining (Stage 1), each training batch consists of 4 labeled samples, while in fine-tuning (Stage 2), each batch includes 2 labeled and 2 unlabeled samples. Throughout training, synthetic prompts are supplied to the frozen generalized teacher to guide supervision, while the student model performs prompt-free inference. For 2D inputs, bounding-box prompts are created by shifting ground-truth annotations or pseudo-label masks by 20 pixels. For 3D inputs, point prompts are sampled either from labeled voxels or from regions with low predictive uncertainty. Following [33], we adopt a Gaussian ramp-up function , where denotes the current iteration and , to progressively increase the supervision strength. The distillation weight is set as with , while the entropy threshold is defined as to gradually relax the filtering criterion [41]. Only voxels with entropy below are supervised, with a fixed loss weight [3].
@@ 4.2.1 Comparison on 2D MRI Segmentation (OASIS)
We evaluate UnCoL on the OASIS dataset across different annotation ratios. As shown in Table 1, UnCoL consistently outperforms existing methods using only 5-10% labeled data. While MedSAM [23], pretrained on brain MRI datasets like FeTA [9], achieves reasonable zero-shot performance (average Dice score of 76.49%), its effectiveness is limited by anatomical and distributional shifts, reflecting the general limitations of zero-shot foundation models under domain shift. UnCoL overcomes this by using uncertainty-guided supervision to adapt effectively with minimal labels.
@@
We observe two major failure modes in prior methods in Fig. 5: (1) TL-FL boundary leakage in the dissection plane due to low contrast, and (2) misclassification of thrombosed FL as background or TL, especially near the entry tear. Clinically, accurate TL-FL separation is essential for assessing blood flow, lumen geometry, and dissection subtype, while FL extent and morphology are closely linked to disease severity and prognosis [11]. TL-FL delineation errors are therefore among the most critical, as they may distort lumen measurements and obscure high-risk regions. UnCoL reduces boundary leakage and better preserves low-contrast, heterogeneous FL regions by integrating global vascular priors with localized refinements, resulting in more clinically consistent TL-FL separation, particularly in high-uncertainty FL areas.
| Method | Model | DPKD | UAPL | OASIS | Pancreas-CT | ImageTBAD | ||||||||||||
| Stu. | G-Tch | S-Tch | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | DSC | Jaccard | 95HD | ASD | ||||
| Baseline | ✓ | 66.395.55 | 56.715.04 | 24.011.55 | 9.843.17 | 56.9616.50 | 41.3813.66 | 18.6912.60 | 5.383.90 | 55.8521.20 | 42.4118.96 | 14.1911.74 | 3.654.56 | |||||
| ✓ | 76.490.67 | 62.680.65 | 7.300.07 | 2.230.04 | 79.374.55 | 66.026.22 | 5.521.79 | 1.550.54 | 29.6119.11 | 19.0414.43 | 13.235.58 | 4.441.94 | ||||||
| ✓ | ✓ | 84.024.15 | 76.413.80 | 12.863.58 | 4.371.72 | 64.2811.62 | 48.4012.10 | 28.1914.54 | 8.925.15 | 64.6715.93 | 50.0415.83 | 7.486.06 | 2.001.43 | |||||
| Data-Efficient Pretraining | ✓ | ✓ | ✓ | 93.652.54 | 88.164.26 | 3.149.13 | 0.901.62 | 73.229.95 | 58.6811.44 | 10.396.57 | 3.381.52 | 52.8316.38 | 41.2716.88 | 4.434.75 | 1.301.27 | |||
| ✓ | ✓ | ✓ | 93.172.17 | 87.293.71 | 2.395.52 | 0.831.28 | 72.669.96 | 57.9811.68 | 11.905.92 | 3.761.60 | 67.6316.48 | 53.3317.06 | 5.444.84 | 1.631.60 | ||||
| ✓ | ✓ | ✓ | ✓ | 93.693.75 | 88.346.03 | 1.950.94 | 0.610.50 | 74.129.49 | 59.7411.09 | 10.796.63 | 3.381.78 | 69.7317.56 | 56.0218.65 | 5.927.43 | 1.641.38 | |||
| Semi-Supervised Fine-tuning | ✓ | ✓ | ✓ | ✓ | ✓ | S-Tch-PL | 93.082.32 | 87.143.91 | 2.445.28 | 0.770.87 | 80.086.11 | 67.188.05 | 7.204.10 | 2.151.08 | 71.261.58 | 57.441.54 | 4.841.75 | 1.380.40 |
| ✓ | ✓ | ✓ | ✓ | ✓ | Avg-Tch-PL | 92.972.32 | 86.953.90 | 2.485.24 | 0.760.77 | 77.297.58 | 63.569.20 | 7.575.13 | 2.321.40 | 70.791.68 | 57.161.22 | 4.871.77 | 1.300.40 | |
| ✓ | ✓ | ✓ | ✓ | ✓ | UAPL | 94.420.29 | 89.510.47 | 2.063.27 | 0.610.63 | 80.405.82 | 67.607.83 | 6.863.96 | 1.850.94 | 70.976.35 | 57.176.81 | 4.832.22 | 1.300.58 | |
@@ 4.3.1 Effect of Teacher Model Configuration
To assess the influence of different teacher configurations, we first compare the baseline model without teacher guidance, single-teacher settings, and the dual-teacher setup. As shown in the top block of Table 4, the Generalized Teacher (G-Tch) increases Dice over the student-only baseline from 66.39% to 76.49% on OASIS and from 56.96% to 79.37% on Pancreas-CT, demonstrating effective prior transfer from the foundation model. However, when encountering unseen anatomy, its performance collapses to 29.61% on ImageTBAD (vs. 55.85% for the student), even with ground-truth-derived prompts, suggesting its domain dependency. The Specialized Teacher (S-Tch), derived via an EMA of the student model, provides stable gains (84.02%, 64.28%, and 64.67%) but underperforms the generalized teacher on well-studied domains (e.g., 64.28% vs. 79.37% on Pancreas-CT). These results show that neither teacher alone is consistently reliable, motivating a dual-teacher design to balance generalization and specialization. Since directly combining two imperfect teachers can propagate errors, DPKD and UAPL are introduced to selectively transfer reliable knowledge, as validated in later ablations.
@@ 4.3.3 Pseudo-Label Quality Analysis
As shown in Table 5, the S-Tch produces more accurate and stable pseudo-labels than the G-Tch, achieving Dice scores of 94.35%, 72.77%, and 66.48% on OASIS, Pancreas-CT, and ImageTBAD, respectively. This improvement is consistent with the effect of transferring generalized semantics, as evidenced by the S-Tch baseline without G-Tch assistance in Table 4 (+10.33%, +8.49%, and +1.61% Dice gains). The G-Tch performs well on OASIS (82.11%) and Pancreas-CT (60.24%) but drops to 15.98% on ImageTBAD, indicating limited domain adaptability at the prediction level. In contrast, the student improves from the specialized teacher to 69.73% Dice on ImageTBAD (+3.25%), showing that representation-level knowledge can still be effectively leveraged, even when G-Tch’s pixel-wise predictions are inaccurate. Overall, these results suggest that our pretraining yields reliable pseudo-labels that adapt well to dataset-specific characteristics, providing a solid foundation for uncertainty-aware fine-tuning.
@@ 4.3.5 Effect of on Model Performance
We evaluate the sensitivity of UnCoL to the pseudo-label balancing coefficient , which controls the trade-off between supervision from ground-truth labels and pseudo-labels during semi-supervised training. Consistent with prior work [3], we set as the default. Figure 7 shows that UnCoL maintains strong performance across a wide range of values. On ImageTBAD, the highest Dice score of 79.24% is obtained at , while Pancreas-CT reaches its peak Dice of 83.91% at . Boundary-based metrics such as the ASD exhibit similar stability. These results suggest that UnCoL is robust to the choice of , with the uncertainty-aware pseudo-labeling mechanism consistently improving performance without the need for precise hyperparameter tuning.
@@ 4.4 Comparison with Classic Fine-tuning Strategies
Experiments in Table 6 follow the fine-tuning paradigm summarized in [12], which provides a unified analysis of classical adaptation strategies for medical foundation models. Our method demonstrates two practical advantages over these tuning baselines. First, as a semi-supervised approach, UnCoL exhibits superior label efficiency. Under the same semi-supervised protocol, it surpasses fine-tuning baselines on both OASIS and ImageTBAD, even when those baselines are trained with partial or full annotations, highlighting its robustness to complex anatomy and severe distribution shifts. On Pancreas-CT, UnCoL achieves clear gains over fine-tuning methods in partially labeled regimes, particularly on region-overlap metrics (Dice and Jaccard), while the boundary-based metrics (95HD and ASD) are comparable to LoRA under partial and full supervision, indicating similar boundary localization accuracy. Second, our method operates in a fully prompt-free manner, whereas the SAM-based fine-tuning baselines require test-time prompts (e.g., bounding boxes or points), introducing additional annotation overhead. Taken together, UnCoL combines strong label efficiency with prompt-free deployment, offering a practical alternative to fine-tuning-based adaptation in semi-supervised medical image segmentation.
@@ 4.5.1 Discriminability
The discriminative power of uncertainty is measured using the area under the ROC curve (AUROC), which quantifies how well uncertainty distinguishes between correct and incorrect predictions. Higher AUROC indicates stronger correlation between uncertainty and voxel-wise error. As shown in Fig. 8(a), UnCoL achieves the highest AUROC on OASIS (0.995) and Pancreas-CT (0.982), and performs competitively on ImageTBAD (0.939). The relatively lower AUROC on ImageTBAD is consistent with domain shifts and complex anatomical variations. Nonetheless, UnCoL still achieves the best Dice and boundary metrics on this dataset (Table 3), indicating that uncertainty estimation remains reliable under challenging conditions. Overall, these results show that our dual-teacher framework improves both segmentation accuracy and uncertainty discriminability.
@@ 4.5.2 Calibration
Reliability diagrams visualize the relationship between predictive confidence and observed accuracy [13]. Perfect calibration corresponds to points lying on the diagonal (confidence = accuracy). As shown in Fig. 8(b), UnCoL exhibits strong calibration across all datasets, with small confidence deviations and low calibration errors measured by Expected Calibration Error (ECE ), Negative Log-Likelihood (NLL ), and Brier score (). These quantitative measures confirm that UnCoL’s predicted confidence values align well with voxel-level correctness.
@@ 4.5.3 Selective prediction reliability
Risk-coverage curves evaluate how uncertainty supports selective prediction by progressively excluding high-uncertainty voxels [8]. A reliable uncertainty measure yields lower selective risk at any target coverage. As shown in Fig. 8(c), UnCoL’s entropy-based uncertainty consistently achieves the lowest risk compared with energy scores, MC Dropout variance, and MC Dropout-based mutual information, indicating that low-uncertainty regions tend to correspond to more reliable predictions. Collectively, these results indicate that UnCoL produces uncertainty estimates that are both discriminative and well-calibrated, effectively supporting pseudo-label selection and consistency regularization under domain-shifted and data-scarce scenarios.